A sugar cone has a slant height of 12 cm and is 4cm across. Find the area needed to make the sugar cone.
Answer by RAY100(1637)

SA of cone = 1/2(2pir)l=pi *r*l = pi * 2 * 12 = 24 pi = 75.4 cm^2